# Aligning Large Language Models with Human Preferences through Representation Engineering

### v1

@@ Abstract
Aligning large language models (LLMs) with human preferences is crucial for enhancing their utility in terms of helpfulness, truthfulness, safety, harmlessness, and interestingness. Existing methods for achieving this alignment often involves employing reinforcement learning from human feedback (RLHF) to fine-tune LLMs based on human labels assessing the relative quality of model responses. Nevertheless, RLHF is susceptible to instability during fine-tuning and presents challenges in implementation. Drawing inspiration from the emerging field of representation engineering (RepE), this study aims to identify relevant representations for high-level human preferences embedded in patterns of activity within an LLM, and achieve precise control of model behavior by transforming its representations. This novel approach, denoted as Representation Alignment from Human Feedback (RAHF), proves to be effective, computationally efficient, and easy to implement. Extensive experiments demonstrate the efficacy of RAHF in not only capturing but also manipulating representations to align with a broad spectrum of human preferences or values, rather than being confined to a singular concept or function (e.g. honesty or bias). RAHF’s versatility in accommodating diverse human preferences shows its potential for advancing LLM performance.
@@ 1 Introduction
While large language models (LLMs) learn broad-ranging world knowledge and a degree of reasoning proficiency, precise control over their behavior proves challenging due to the unsupervised nature of their pre-training Radford et al. (2018, 2019); Brown et al. (2020); Bubeck et al. (2023); Touvron et al. (2023). For each query, instruction-tuned LLMs Wei et al. (2021); Chung et al. (2022); Touvron et al. (2023) exhibit the capacity to generate multiple responses that are both semantically and syntactically coherent by some sampling techniques. While such ability enables the models to provide diversity that is essential for chat agents, some responses may contain harmful, unethical, socially biased and negative, even illegal content Srivastava et al. (2022); Thoppilan et al. (2022); Bubeck et al. (2023); Wang et al. (2023). Although many responses remain within acceptable bounds, preferences may exist among them, leading to a desire to prioritize specific outputs over others.
@@
In addressing the aforementioned challenges posed by RL-based methods, several computationally lightweight alternatives have been proposed to simplify the human preference matching process. Two prominent paradigms among these alternatives include contrastive learning Rafailov et al. (2023); Zhao et al. (2023); Yuan et al. (2023) and Hindsight instruction relabeling (HIR) Zhang et al. (2023); Liu et al. (2023). Contrastive learning-based methods optimize a language model policy by increasing the relative probability of preferred responses over dispreferred ones, while HIR methods transforms human feedback into instructions by relabeling the original ones, indicating the relative quality of provided responses. A common characteristic shared by these two paradigms is their capability to align language models with human preferences through reward-free fine-tuning.
@@
However, the reward-free fine-tuning is vulnerable to the presence of noisy data or incorrect labels in a training set comprising a collection of preference-annotated response pairs. Instances of dull sentences or very brief responses may appear repeatedly in such a training set, potentially introducing bias into the models. The exclusion of such instances from the training set renders it impossible for LLMs to glean insights of human preferences expressed in these instances. In contrast, RL-based methods adopt a different strategy, wherein a reward function is first extracted from a dataset of response rankings, and then this reward function can be applied to train a LLM, effectively mitigating the model’s direct exposure to noisy data or incorrect labels within the dataset.
@@
In this study, we aim to seek for a computationally lighter and reward-free algorithm that can effectively harness human preference expressed in datasets meanwhile safeguarding LLMs from the influence of noisy data. Inspired by the recent advance in representation engineering Zou et al. (2023), we initially locate relevant representations and activity patterns associated with high-level human preferences within an LLM, and subsequently gain precise control over its behavior by manipulating its internal representations. In the neural architecture, network weights determine neural activity, neural activity determines the networks’ output, and the networks’ output determines the networks’ behavior.n Instead of focusing on neurons and their connections we see aligning LLMs with human feedback as an outcome of representational spaces, implemented by patterns of activity across populations of neurons. We first identify the differences in model activities between preferred and dispreferred stimuli, and then control over its behavior by leveraging the identified differences in representations (see Figure 1). We introduce two methods for controlling representations and demonstrate the efficacy of these representation engineering (RepE) approaches in aligning LLMs with a broad spectrum of human preferences through a collection of response pairs.
@@
To validate the effectiveness of our approach in aligning with human preferences, we conducted extensive comparative experiments on the generated results. Our method outperformed other RL-free approaches in human evaluations and automated metrics such as reward model scores, GPT-4 evaluations, and also achieved results comparable to RLHF. Notably, the underlying algorithms exhibit simplicity in implementation and straightforwardness in training.
@@ 2 Related Work
Extensive research has devoted into the exploration of RL from human feedback through ratings or rankings, spanning tasks from NL-to-SQL conversion Zhong et al. (2017), machine translation Kreutzer et al. (2018), task-oriented dialogue systems Su et al. (2019); Zhang et al. (2019); Takanobu et al. (2019), summarization Stiennon et al. (2020), story-telling Ziegler et al. (2019) to instruction-following Ouyang et al. (2022); Ramamurthy et al. (2023). Typically, these methods involve the fitting of a reward model to a dataset of human preferences, followed by the optimization of a LLM policy to generate responses with high reward, using RL algorithms such as REINFORCE Williams (1992) or proximal policy optimization Schulman et al. (2017). Despite the attractiveness of leveraging human preferences that are easier to collect than expert demonstrations, training LLMs with RL poses significant practical challenges, which is attributed to the sensitivity of RL to hyperparameters and the inherent instability during training.
@@
To enhance transparency and controllability of neural networks, Zou et al. (2023) introduced representation engineering (RepE) as a methodology, drawing an analogy between understanding deep neural networks through representation tomography and studying brains via neuroimaging techniques. Their work demonstrated the efficacy of RepE in addressing diverse safety-related challenges such as truthfulness, honesty, and hallucination. This study falls in line with recent research findings and extends its application to aligning LLMs with a wide spectrum of human preferences. Our study introduces two novel methods to instruct LLMs on human preferences first, and then extract differences in model activities between preferred and dispreferred stimuli. These differences in activity patterns serve as a foundation for manipulating the model’s behavior, leading to the generation of responses that better align with human preferences. Due to the lightweight computational advantages of parameter-efficient fine-tuning techniquesHoulsby et al. (2019); Lester et al. (2021); Hu et al. (2021); Wu et al. (2023) , these techniques are utilized to fit the disparity in activity patterns. In contrast to the approach adopted by Zou et al. (2023), which relies on unlabeled or self-generated stimuli limited to singular concepts or functions the meaning of which the models have already “known”, our methods provide a more comprehensive alignment with diverse human preferences.
Figure 2: The procedure of RAHF. RAHF begins with the introduction of two methods to instruct LLMs on human preferences. One approach involves training a single LLM to discern the relative quality of responses (RAHF-SCIT), while the other employs dual LLMs to model preferred and dispreferred responses separately (RAHF-DualLLMs). Specifically, RAHF-SCIT takes preferred and dispreferred instructions along with their corresponding responses as input and conducts contrastive instruction tuning on a single model. RAHF-DualLLMs, on the other hand, performs supervised training by taking preferred and dispreferred responses into different models. Subsequently, we obtain activity patterns by stimulating the model with different prompts. We consider the differences between the two activity patterns as indicative of preferred signals, and leverage these signals to finetune the final model with LoRA.
@@ 3 Method
We begin by instructing LLMs on human preferences with a set of preference-annotated response pairs. Secondly, we collect the activity patterns of LLMs when exposed to stimuli that are preferred or dispreferred. The differences in these patterns serve as the foundation for manipulating LLMs, enabling them to generate responses more closely aligned with human values. We introduce two novel methods for instructing LLMs on human preferences and extracting their activity patterns: one involving a single LLM (trained to discern the relative quality of responses) and the other employing dual LLMs (“a good guy and a bad guy”). Finally, we construct the final model by training a low-rank adapterHu et al. (2021) to fit the disparity in activity patterns.
@@ 3.1 Instructing LLMs on Human Preferences
Zou et al. (2023) validated the effectiveness of activity patterns extracted from alignment fine-tuned models, such as LLaMA-2-chat, in capturing concepts like truthfulness and honesty. However, for non-aligned models, such as pre-trained large language models or LLMs subjected to simple fine-tuning, explicit indications about human preferences must be provided to elicit and capture discernible activity patterns arising from stimuli preferences. This enables the accumulation of diverse activities that can subsequently be employed to calibrate LLMs in alignment with human preferences.
To attain this objective, we explored two different methods. The first involves instructing a single LLM using preference-annotated response pairs, employing contrasting instructions like Hindsight Zhang et al. (2023). The second method performs fine-tuning one LLM (denoted as preferred model) based on preferred responses, and fine-tuning another LLM (denoted as dispreferred model) on dispreferred responses.
3.1.1 Instruction with a Single Model
Within the proposed framework, the Single LLM Method focuses on fine-tuning a Single Large Language Model through Contrastive Instruction Tuning (SCIT). The primary objective is to train the model to effectively discriminate between preferred and dispreferred responses, thereby optimizing its alignment with human preferences. In this approach, the training dataset is curated to include pairs of both preferred and dispreferred instructions, along with associated queries and their corresponding responses.
@@ 3.1.1 Instruction with a Single Model
Inspired by HIR Zhang et al. (2023), for instructions linked to positive preferences, the goal is to heighten the probability of generating preferred responses while concurrently diminishing the probability of generating dispreferred responses. Conversely, for instructions associated with negative preferences, the objective is to elevate the probability of generating dispreferred responses and reduce the probability of generating preferred responses.
@@
Formally, let $D$ represent the training dataset, with $q_{i}$ denoting the query, $r_{i}$ representing the response, and $p_{i}$ indicating the instruction (positive or negative). The fine-tuning of the LLM involves maximizing the following objectives:
$\pi(\theta^{*})$ $=\arg\max_{\theta}\sum_{(p_{i},q_{i},r_{i})\in D}\log\frac{\exp{(P^{+})}}{\exp{(P^{+})}+\exp{(P^{-})}}$ (1)
@@
where $P^{+}$$=\pi(r_{i}\mid p_{i},q_{i};\theta)$, $P^{-}$$=\pi(r_{i}$$\mid p_{i}^{*},q_{i};\theta)$ and $p_{i}^{*}$ denotes the opposite instruction, ensuring a contrast between preferred and dispreferred cases.
Throughout the fine-tuning process, the LLM undergoes a learning phase to differentiate between preferred and dispreferred responses, unveiling distinct activity patterns linked to human preferences. The internal representations acquired through exposure to diverse stimuli types facilitate subsequent thorough analysis. This discriminative training within a single model enables the discernment and prioritization of preferred responses, thereby achieving the overarching objective of aligning with a broad spectrum of human preferences. Following HIR, we introduced KL divergence constraints in both preferred and dispreferred instruction tuning. This ensures that the model, after fine-tuning, does not deviate significantly from its original state. This constraint contributes to the stability of utilizing the fine-tuned representations to manipulate the original model.
@@ 3.1.2 Preference Instruction with Dual Models
In Dual LLMs method, our aim is to train two LLMs with distinct tendencies: one model is inclined to generate preferred responses, while the other tends to produce dispreferred responses. To achieve this objective, we employ paired preference data to conduct supervised fine-tuning of the LLMs. Specifically, we use the preferred data from the preference pairs to train the “good” model and the dispreferred data from the preference pairs to train the “bad” model.
@@
Through this training process, the “good” model and “bad” model have respectively learned the activity patterns associated with human-preferred and dispreferred responses.
@@ 3.2 Collecting Activity Patterns
In the process of extracting activity patterns, we utilize stimulus pairs <$p^{+},q,r$> and <$p^{-},q,r$> to elicit representations from the intermediate layers of the model. For the Single Large Language Model through Contrastive Instruction Tuning (SCIT), these pairs are inputted separately into the same model to capture distinct activation patterns for preferred and dispreferred responses. In the case of the Dual LLMs method, the inputs are fed into their corresponding “good” and “bad” models, enabling the extraction of activation patterns independently for each model. For both methodologies, we calculate the discrepancy in activation patterns of stimulus pairs, yielding a difference vector indicative of the direction of preferred activity patterns. Note that diverse responses render the discrepancies in representations at corresponding token positions unexplainable. Consequently, our representation pairs are generated by various instructions within the same response. Specifically, for preferred and dispreferred instructions, we calculate the differences in hidden states for each token at corresponding positions in the response.
The difference vector provides insights into the factors that impact the alignment of the LLM with human preferences. Subsequently, we perturb the model’s original representation by incorporating the difference vectors. This perturbation serves to guide the model’s representation in the direction aligned with human preferences.

### v3

@@ Abstract
Aligning large language models (LLMs) with human preferences is crucial for enhancing their utility in terms of helpfulness, truthfulness, safety, harmlessness, and interestingness. Existing methods for achieving this alignment often involve employing reinforcement learning from human feedback (RLHF) to fine-tune LLMs based on human labels assessing the relative quality of model responses. Nevertheless, RLHF is susceptible to instability during fine-tuning and presents challenges in implementation. Drawing inspiration from the emerging field of representation engineering (RepE), this study aims to identify relevant representations for high-level human preferences embedded in patterns of activity within an LLM and achieve precise control of model behavior by transforming its representations. This novel approach, denoted as Representation Alignment from Human Feedback (RAHF), proves to be effective, computationally efficient, and easy to implement. Extensive experiments demonstrate the efficacy of RAHF in not only capturing but also manipulating representations to align with a broad spectrum of human preferences or values, rather than being confined to a singular concept or function (e.g. honesty or bias). RAHF’s versatility in accommodating diverse human preferences shows its potential for advancing LLM performance. Code is available at https://github.com/LiuAmber/RAHF.
Aligning Large Language Models with Human Preferences
through Representation Engineering
Wenhao Liuthanks:   These authors contributed equally., Xiaohua Wang11footnotemark: 1, Muling Wu, Tianlong Li, Changze Lv Zixuan Ling, Jianhao Zhu, Cenyuan Zhang, Xiaoqing Zhengthanks:   Corresponding author., Xuanjing Huang School of Computer Science, Fudan University, Shanghai, China {whliu22,xiaohuawang22}@m.fudan.edu.cn {zhengxq,xjhuang}@fudan.edu.cn
@@ 1 Introduction
While large language models (LLMs) learn broad-ranging world knowledge and a degree of reasoning proficiency, precise control over their behavior proves challenging due to the unsupervised nature of their pre-training Radford et al. (2018, 2019); Brown et al. (2020); Bubeck et al. (2023); Touvron et al. (2023). For each query, instruction-tuned LLMs Wei et al. (2021); Chung et al. (2022); Touvron et al. (2023) exhibit the capacity to generate multiple responses that are both semantically and syntactically coherent by some sampling techniques. While such ability enables the models to provide diversity that is essential for chat agents, some responses may contain harmful, unethical, socially biased, and negative, even illegal content Srivastava et al. (2022); Thoppilan et al. (2022); Bubeck et al. (2023); Wang et al. (2023).
@@
In addressing the aforementioned challenges posed by RL-based methods, several computationally lightweight alternatives have been proposed to simplify the human preference-matching process. Two prominent paradigms among these alternatives include contrastive learning Rafailov et al. (2023); Zhao et al. (2023); Yuan et al. (2023) and Hindsight instruction relabeling (HIR) Zhang et al. (2023); Liu et al. (2023). Contrastive learning-based methods optimize a language model policy by increasing the relative probability of preferred responses over dispreferred ones, while HIR methods transform human feedback into instructions by relabeling the original ones, indicating the relative quality of provided responses. A common characteristic shared by these two paradigms is their capability to align language models with human preferences through reward-free fine-tuning.
@@
However, the reward-free fine-tuning is vulnerable to the presence of noisy data or incorrect labels in a training set comprising a collection of preference-annotated response pairs Li et al. (2023b); Dumoulin et al. (2023). Instances of dull sentences or very brief responses may appear repeatedly in such a training set, potentially introducing bias into the models. The exclusion of such instances from the training set renders it impossible for LLMs to glean insights into human preferences expressed in these instances. In contrast, RL-based methods adopt a different strategy, wherein a reward function is first extracted from a dataset of response rankings, and then this reward function can be applied to train an LLM, effectively mitigating the model’s direct exposure to noisy data or incorrect labels within the dataset.
@@
In this study, we aim to seek for a computationally lighter and reward-free algorithm that can effectively harness human preference expressed in datasets meanwhile safeguarding LLMs from the influence of noisy data. Inspired by the recent advance in representation engineering Zou et al. (2023), we initially locate relevant representations and activity patterns associated with high-level human preferences within an LLM, and subsequently gain precise control over its behavior by manipulating its internal representations. In the neural architecture, network weights determine neural activity, neural activity determines the networks’ output, and the networks’ output determines the networks’ behavior. Instead of focusing on neurons and their connections, we see aligning LLMs with human feedback as an outcome of representational spaces, implemented by patterns of activity across populations of neurons. We first identify the differences in model activities between preferred and dispreferred stimuli, and then control its behavior by leveraging the identified differences in representations (see Figure 1). We introduce two methods for controlling representations and demonstrate the efficacy of these representation engineering (RepE) approaches in aligning LLMs with a broad spectrum of human preferences through a collection of response pairs.
@@
To validate the effectiveness of our approach in aligning with human preferences, we conducted extensive comparative experiments on the generated results. Our method outperformed RLHF and other RL-free approaches in human evaluations and automated metrics such as general abilities and GPT-4 evaluations. Notably, the underlying algorithms exhibit simplicity in implementation and straightforwardness in training.
@@ 2 Related Work
Extensive research has been devoted to the exploration of RL from human feedback through ratings or rankings, spanning tasks from NL-to-SQL conversion Zhong et al. (2017), machine translation Kreutzer et al. (2018), task-oriented dialogue systems Su et al. (2019); Zhang et al. (2019); Takanobu et al. (2019), summarization Stiennon et al. (2020), story-telling Ziegler et al. (2019) to instruction-following Ouyang et al. (2022); Ramamurthy et al. (2023). Typically, these methods involve the fitting of a reward model to a dataset of human preferences, followed by the optimization of a LLM policy to generate responses with high reward, using RL algorithms such as REINFORCE Williams (1992) or proximal policy optimization Schulman et al. (2017). Despite the attractiveness of leveraging human preferences that are easier to collect than expert demonstrations, training LLMs with RL poses significant practical challenges, which is attributed to the sensitivity of RL to hyperparameters and the inherent instability during training.
@@
To enhance transparency and controllability of neural networks, Zou et al. (2023) introduced representation engineering (RepE) as a methodology, drawing an analogy between understanding deep neural networks through representation tomography and studying brains via neuroimaging techniques. Their work demonstrated the efficacy of RepE in addressing diverse safety-related challenges such as truthfulness, honesty, and hallucination. This study falls in line with recent research findings and extends its application to aligning LLMs with a wide spectrum of human preferences. Our study introduces two novel methods to instruct LLMs on human preferences first, and then extract differences in model activities between preferred and dispreferred stimuli. These differences in activity patterns serve as a foundation for manipulating the model’s behavior, leading to the generation of responses that better align with human preferences. Due to the lightweight computational advantages of parameter-efficient fine-tuning techniquesHoulsby et al. (2019); Lester et al. (2021); Hu et al. (2021); Wu et al. (2023, 2024), these techniques are utilized to fit the disparity in activity patterns. In contrast to the approach adopted by Zou et al. (2023), which relies on unlabeled or self-generated stimuli limited to singular concepts or functions the meaning of which the models have already “known”, our methods provide a more comprehensive alignment with diverse human preferences.
Figure 2: The procedure of RAHF. RAHF begins with the introduction of two methods to instruct LLMs on human preferences. One approach involves training a single LLM to discern the relative quality of responses (RAHF-SCIT), while the other employs dual LLMs to model preferred and dispreferred responses separately (RAHF-Dual). Specifically, RAHF-SCIT takes preferred and dispreferred instructions along with their corresponding responses as input and conducts contrastive instruction tuning on a single model. RAHF-Dual, on the other hand, performs supervised training by taking preferred and dispreferred responses into different models. Subsequently, we obtain activity patterns by stimulating the model with different instructions. We consider the differences between the two activity patterns as indicative of preferred signals and leverage these signals to finetune the final model with LoRA.
@@ 3 Method
We begin by instructing LLMs on human preferences with a set of preference-annotated response pairs. We introduce two novel methods for instructing LLMs on human preferences and extracting their activity patterns: one involving a single LLM (trained to discern the relative quality of responses) and the other employing dual LLMs (“a good guy and a bad guy”). Secondly, we collect the activity patterns of LLMs when exposed to stimuli that are preferred or dispreferred. The differences in these patterns serve as the foundation for manipulating LLMs, enabling them to generate responses more closely aligned with human values. Finally, we construct the final model by training a low-rank adapterHu et al. (2021) to fit the disparity in activity patterns.
@@ 3.1 Instructing LLMs on Human Preferences
To extract activity patterns from the model that align with human preferences, it is crucial for the model to possess a correct understanding and awareness of these preferences. The effectiveness of extracting activity patterns from alignment fine-tuned models, such as LLaMA-2-chat, in capturing concepts like truthfulness and honesty has been validated by Zou et al. (2023). However, for non-aligned models, such as pre-trained large language models or LLMs subjected to simple fine-tuning, explicit indications of human preferences should be provided to elicit and capture activity patterns induced by stimulus preferences. This capability enables the accumulation of diverse activities, subsequently utilized to calibrate LLMs based on human preferences.
For instructing LLMs on human preferences, we rely on a dataset annotated with human preferences. As mentioned earlier, we employ two methods to achieve this goal. The first method utilizes HindsightZhang et al. (2023), using contrastive instructions to instruct a single LLM. The second method involves fine-tuning two LLMs separately: one (referred to as the preferred model) is fine-tuned based on preferred responses, while the other (referred to as the dispreferred model) is fine-tuned on dispreferred responses.
3.1.1 Preference Instruction with a Single Model
Within the proposed framework, the Single LLM Method focuses on fine-tuning a Single Large Language Model through Contrastive Instruction Tuning (SCIT). This process involves two instructions: one instructs the model to generate responses preferred by humans, while the other guides the model to generate responses dispreferred by humans. Following such fine-tuning, we can optimize the model for consistency with human preferences. We can also stimulate the model to elicit distinct activity patterns by employing different instructions subsequently.
@@ 3.1.1 Preference Instruction with a Single Model
Specifically, the training dataset is curated to include pairs of both preferred and dispreferred instructions, alongside associated queries and their corresponding responses (details on preferred instructions can be found in Appendix A.1). Following HIRZhang et al. (2023), for instructions linked to positive preferences, the fine-tuning objective aims to increase the probability of generating preferred responses while concurrently decreasing the probability of generating dispreferred responses. Conversely, for instructions associated with negative preferences, the objective is to elevate the probability of generating dispreferred responses and reduce the probability of generating preferred responses.
@@
Formally, let $D$ represent the training dataset, with $q_{i}$ denoting the query, $r_{i}$ representing the response, and $p_{i}$ indicating the instruction (positive or negative). The fine-tuning of the LLM involves minimizing the following loss:
$\mathcal{L}$ $=-\sum_{(p_{i},q_{i},r_{i})\in D}(P^{+}+\log\frac{\exp{(P^{+})}}{\exp{(P^{+})}+\exp{(P^{-})}})$ (1)
@@
where $P^{+}$$=\log\pi(r_{i}\mid p_{i},q_{i};\theta)$, $P^{-}$$=\log\pi(r_{i}$$\mid p_{i}^{*},q_{i};\theta)$ and $p_{i}^{*}$ denotes the opposite instruction, ensuring a contrast between preferred and dispreferred cases.
Throughout the entire fine-tuning process, the LLM undergoes a learning phase to distinguish between preferred and non-preferred responses, revealing distinct activity patterns associated with human preferences. Subsequently, these two instructions will serve as stimuli to acquire the model’s internal representations, which will be used for further alignment. This contrastive training relying on preference data enables the achievement of the overarching goal of consistency with a broad spectrum of human preferences, rather than a singular concept.
@@ 3.1.2 Preference Instruction with Dual Models
In the Dual LLMs method, we aim to train two LLMs with distinct tendencies: one model is inclined to generate preferred responses, while the other tends to produce dispreferred responses. To achieve this objective, we employ paired preference data to conduct supervised fine-tuning of the LLMs. Specifically, we use the preferred data from the preference pairs to train the preferred model and the dispreferred data from the preference pairs to train the dispreferred model.
@@
Through this training process, the preferred model and dispreferred model have respectively learned the activity patterns associated with human-preferred and dispreferred responses. Due to the human preference learning conducted in two distinct models, in contrast to SCIT, the Dual LLMs method does not require additional distinct instructions during fine-tuning. Instead, guidance for the model is provided solely through different responses.
@@ 3.2 Collecting Activity Patterns
Following the establishment of comprehension of human preferences by LLMs, we are able to extract representations of what humans prefer and disprefer. Due to the characteristics of autoregressive Transformer language models, the attention mechanism results in tokens at different positions exhibiting distinct representations. The activation representation of a token at the current position is influenced by preceding tokens. Therefore, for a specific pair of query $q$ and response $r$, this pair is concatenated with two instructions from Section 3.1, which guide the model in forming the concept of human preferences and inputted into the model to obtain the intermediate layer hidden states at each position as internal representations.
Figure 3: Examples of Collecting Activity Patterns. To ensure the correspondence between the positions of preferred and dispreferred instructions during the extraction of difference vectors, instruction $p$ and query $q$ are left-padded to the maximum prompt length, while the response $r$ is right-padded to the maximum response length.
Formally, for a given instruction $p$, a decoder model $\pi$, we collect the $l$-th layer’s hidden states of each token for the query-response pair ($q_{i},r_{i}$) within dataset $D$. This can be formalized as follows:
$A_{p,\pi,l}={\pi_{l}(p,q_{i},r_{i})\mid(q_{i},r_{i})\in D}$ (4)
Here, $\pi_{l}$ represents the hidden states output by the neural network’s $l^{th}$ layer. We directly extract the hidden layer states from the neural network as representations. To address the issue of varying response lengths during the activity pattern collection, we concatenated the same response with different instructions as input to ensure the representations were extracted with the same length. Different instructions will elicit distinct activity patterns even though the same response was provided and the differences in the elicited activity patterns can be used to capture the behavior of the models. Such differences can be conceptualized and modeled as the probability of generating the same response conditioned on different instructions. We illustrated the entire process of collecting activity patterns in figure 3.
We obtain the final difference vector by subtracting the hidden states of dispreferred outputs from those of preferred outputs, as expressed by the equation:
$v_{l}=A_{p^{+},\pi,l}-A_{p^{-},\pi,l}$ (5)
This difference vector $v_{l}$ represents the difference in activation patterns produced under the two different stimulus conditions $p^{+}$ and $p^{-}$. Subsequently, we perturb the model’s original representation by incorporating the difference vectors. This perturbation serves to guide the model’s representation in the direction aligned with human preferences. It is noteworthy that, for the Single Large Language Model through Contrastive Instruction Tuning (SCIT), both $A_{p^{+},\pi,l}$ and $A_{p^{-},\pi,l}$ are generated by the same model. In the dual LLMs approach, pairs concatenated with different instructions are inputted into the respective preferred and dispreferred models, thereby enabling the independent extraction of activation patterns from each model.